OpenStudy (anonymous):

The sum of two numbers is 70. The larger number is 4 more than twice the smaller number. What is the smaller number?

OpenStudy (conqueror):

x + y = 70 x = 2y+4

OpenStudy (texaschic101):

smaller number = y larger number = x x + y = 70 x = 2y + 4 now sub in 2y + 4 in for x in the other equation 2y + 4 + y = 70 3y + 4 = 70 3y = 70 - 4 3y = 66 y = 22 x + y = 70 x + 22 = 70 x = 70 - 22 x = 48 check.. x = 2y + 4 48 = 2(22) + 4 48 = 44 + 4 48 = 48 (correct) larger number (x) = 48 smaller number (y) = 22

OpenStudy (anonymous):

Thanks.

OpenStudy (texaschic101):

anytime :)